Condense the expression
Ln 20+2ln 1/2 - ln x

Answer by mamiya(56)

Condense the expression
Ln 20+2ln 1/2 - ln x
first, we have to remember that exponential logarithm turns a division into a subtraction and a product into an addition and an exponent into a product, and vice-verca.
with that in mind, we have ,
ln 20 + 2ln1/2 - ln x = ln 20 - ln x + ln(1/2)^2
= ln 20/x + ln 1/4
= ln (20/x)(1/4)
= ln (20/4x)
= ln 5/x
so the condense form of ln 20 + 2ln1/2 - ln x is ln 5/x